OpenStudy (anonymous):

defferentiat the following: 5*e^x*sin(x)

OpenStudy (anonymous):

d/dx [ 5*e^x*sin(x) ] = 5 * d/dx [ e^x * sin(x) ] = chain rule = 5 * [ (d (e^x)/dx) * sin(x) + e^(x) * (d(sin(x))/dx) ] = 5 * [e^x * sin(x) + e^x * cos(x) ] = 5*e^(x)*[cos(x) + sin(x)]

OpenStudy (anonymous):

btw understand the question with the ladder ?

OpenStudy (anonymous):

yes i got the one with the ladder. Thanks so much Also thanks for explaining this question too.